Through the two ends of a glass tube of length 200 cm HCl gas and NH3 are allowed to enter. At what distance the white ring of ammonium chloride will first appear from HCl end

Grade:12

1 Answers

Vikas TU
7 years ago
Let x be the length from the HCl end and therfore, (2-x) from the other end.
for x length rate of diffusion that is,
dx/dt = PA/(root(36.5T))
d(2-x)/dt = PA/root(17T)
dividing both we get,
x/(2-x) = > root(17/36.5) 
x = 2/2.466 = > 0.81 m or 81 cm.